# ProTrix: Building Models for Planning and Reasoning over Tables with Sentence Context

### v2

@@ 3.3 Free-form Answer Tasks
We run GPT-3.5-turbo and TableLlama Zhang et al. (2023a) as our baselines. The prompt for each model is shown in Table 15 in Appendix. We also use the result of fine-tuning method using T5-large, and human performance from Nan et al. (2022) as baselines. Notably, the results from Nan et al. (2022) are evaluated with the original setting where the highlighted cells are provided as input instead of the whole table.
@@
Since the response of our model contains step-by-step reasoning over symbolic code and natural language, BLEU Papineni et al. (2002) would underestimate the performance of our model. BLEU also can not be used to evaluate the correctness and faithfulness of the responses. We sample 100 cases from the dataset to perform human evaluation following Nan et al. (2022). The evaluation is based on four criteria: (1) fluency if an answer is natural and grammatical; (2) correctness if an answer is correct; (3) adequacy if an answer contains all the information that is asked by the question; (4) faithfulness if an answer is faithful and grounded to the contents of the table.
@@ Results
4 Ablation study
To better evaluate the effectiveness of fine-tuning with our Plan-then-Reason framework, we experiment with 3 other fine-tuning frameworks based on TrixInstruct. (1) w/o Planning: We split each instance in TrixInstruct into planning and reasoning parts. We train our model with only the reasoning part of the training instances. This can be considered as distilling the reasoning pattern of TabSQLify Nahid and Rafiei (2024). (2) w/o Reasoning: Similar to (1), we finetune the model with only the planning part of the training instances. (3) w/o Planning and Reasoning: We finetune the model to generate answers directly. This is similar to the end-to-end QA paradigm.
@@ 4 Ablation study
5 Related Work
Prompting Methods for LLMs
Large language models can be guided to solve tasks in a step-by-step manner Wei et al. (2022); Hao et al. (2023). Chen (2023) first utilizes Chain-of-Thought Wei et al. (2022) to enhance the reasoning of LLMs on tabular tasks and points out that textual reasoning can not generalize to large tables directly. Researchers prompt the model to select relevant rows and columns as one step in the chain of reasoning to enable LLM to focus on the following reasoning step Jiang et al. (2023); Ye et al. (2023); Wang et al. (2024). Chen et al. (2022) proposes Program-of-Thoughts (PoT) that answers a question in programming language. Compared with textual reasoning, program-based reasoning is executed by a code interpreter, achieving high-precision reasoning in complex tabular or mathematical questions. Binder Cheng et al. (2022b) binds LLMs as API calls within a Python or SQL program to address the program-unsolvable aspect of the queries. Liu et al. (2023b) proposes mix self-consistency that combines the potential of both textual and program-based reasoning. Researchers have also attempted ReAct Yao et al. (2023) style prompting for tabular tasks. ReAcTable Zhang et al. (2023b) reasons with SQL or Python tools in multiple turns. Chain-of-Table Wang et al. (2024) formats table reasoning as specific table operations. ReAct-style prompting mainly focuses on the next acting step based on feedback from the last action without considering the whole reasoning chain. Our Plan-then-Reason framework can plan the whole reasoning chain before acting and answer questions more efficiently and accurately (see efficiency analysis in Appendix B) while blending the advantages of textual and program-based reasoning methods.
Finetuned Models
Various pre-trained models are proposed for tabular tasks Yin et al. (2020); Wang et al. (2021); Iida et al. (2021); Deng et al. (2022); Yang et al. (2022); Jiang et al. (2022); Liu et al. (2021). But they often are limited to one specific downstream fine-tuning task. As for models with generalizability, Liu et al. (2023a) mix symbolic SQL execution task with FLAN task to further fine-tune FLAN-T5 to improve zero-shot tabular question answering performance. Li et al. (2023) fine-tunes models with a large synthesized dataset of table manipulation and data augmentation to serve as a table-foundation model that understands table structures. TableLlama Zhang et al. (2023a) and StructLM Zhuang et al. (2024) collect an instructing tuning set that covers diverse tables and tasks and finetune Llama to obtain a generalist model without table pretraining. Compared with existing generalist models that are expected to generate answers directly, ProTrix is interpretable by generating accurate and faithful explanations.
6 Conclusions
In this paper, we propose Plan-then-Reason framework to answer different types of user queries over tables with sentence context. It understands the commonsense and concepts in the query and plans the reasoning steps over programs and languages. We construct an instruction tuning set TrixInstruct to finetune models to obtain such planning and reasoning abilities with only 6k examples. The experiments show that our resulting models ProTrix and ProTrix-Coder generalize to unseen tabular tasks with sentence context and produce accurate and faithful explanations. Our work highlights the required abilities for generalist models over tabular tasks with sentence context, and paves the way for future research directions.
Limitations
Appendix A Implementation Details
We fully fine-tune Llama-2 7B Touvron et al. (2023) and CodeLlama-7B Roziere et al. (2023) with our instruction tuning set with the context of length 4096. We set the learning rate as 5e-6 and the batch size as 32. The training process uses a cosine scheduler with a 3% period for 3 epochs. We utilize DeepSpeed training with ZeRO-3 stage Rasley et al. (2020). Our model is trained with 4 NVIDIA A40 GPUs (48GB) and the whole training process takes about 5 hours.
For finetuned SOTA methods in Table 1, we report the performance of OmniTab Jiang et al. (2022) for WikiTQ, TAPEX Liu et al. (2021) for WikiSQL, PASTA Gu et al. (2022) for TabFact, finetuned BERT for SciTAB Lu et al. (2023), S3HQA Lei et al. (2023) for HybridQA and APOLLO Sun et al. (2022) for TATQA. For FEVEROUS, we run DCUF Hu et al. (2022) on our training and development set of FEVEROUS and obtain an accuracy of 75.9%. Notice that S3HQA uses a more precise sentence retriever compared to ours and DCUF leverages an additional retriever to select 25 table cells as input.
Appendix B Efficiency Analysis
@@
In Table 4, we compare the number of annotated in-context demonstrations and average API calls of different prompting methods on WikiTQ. It shows that our Plan-then-Reason framework can effectively reach the final answer with less in-context demonstrations and API calls. Binder Cheng et al. (2022b), Dater Ye et al. (2023), Chain-of-Table Wang et al. (2024) and TabSQLify Nahid and Rafiei (2024) requires more than 10 annotated examples for inference while our Plan-then-Reason framework only needs one in-context demonstration. Binder, Dater and MIX SC predicts the answer with self-consistency mechanism which needs at least 10 API calls. Chain-of-Table follows the style of ReAct Yao et al. (2023). The average API calls to reach the final answer is 10.1 while Plan-then-Reason only requires 2 API calls on average.
Appendix C Training Dataset Analysis
@@
Our instruction tuning dataset is extracted from GPT-4 responses. We filter out the responses that have inconsistent final answers with the original dataset annotations, but the reasoning process of the responses in the training set has not been fully validated. We perform a quality analysis of our training dataset. We categorize the cases into the following types: (1) Planning error: the response fails to generate a correct plan to answer the question, (2) SQL error: The response generates SQL containing columns that do not exist at all or the SQL simply can not be executed. (3) Execution error: the execution result given by GPT-4 is wrong and (4) Reasoning error: the model achieves a correct answer with wrong reasoning.We demonstrate examples for each error type in Table 8. We sample 50 cases with sentence context and 50 cases without sentence context from TrixInstruct to perform manual evaluation.
@@
Appendix D Analysis of Breakdown Performance
D.1 Program-Unsolvable Queries
@@
From Table 6, we can observe that ProTrix-Coder achieves the highest accuracy on program-unsolvable queries compared with fine-tuned methods. It suggests TrixInstruct can teach a model to understand commonsense and general concepts in the query and adaptatively plan to reason with programs or languages. ProTrix-Coder still falls behind TAPEX and TaCube on the program-solvable subset. But these models require table pretraining which is computationally expensive. ProTrix-Coder surpasses the previous generalist model by 1.5% and 4.7% on program unsolvable and solvable subsets, indicating the effectiveness of the proposed Plan-then-Reason framework.
@@
D.2 Combining Tables and Sentences
@@
Appendix E More Results
@@
Appendix F Case Study

### v3

@@ 3.3 Free-form Answer Tasks
We run GPT-3.5-turbo and TableLlama Zhang et al. (2023a) as our baselines. The prompt for each model is shown in Table 16 in Appendix. We also use the result of finetuning method using T5-large, and human performance from Nan et al. (2022) as baselines. Notably, the results from Nan et al. (2022) are evaluated with the original setting where the highlighted cells are provided as input instead of the whole table.
@@
Since the response of our model contains step-by-step reasoning over symbolic code and natural language, BLEU Papineni et al. (2002) would underestimate the performance of our model. BLEU also can not be used to evaluate the correctness and faithfulness of the responses. We sample 100 cases from the dataset to perform human evaluation following Nan et al. (2022). The evaluation is based on four criteria: (1) fluency: if an answer is natural and grammatical; (2) correctness: if an answer is correct; (3) adequacy: if an answer contains all the information asked by the question; (4) faithfulness: if an answer is faithful and grounded to the contents of the table.
@@ Results
4 Analysis
4.1 Ablation study
To better evaluate the effectiveness of finetuning with our Plan-then-Reason framework, we experiment with 3 other finetuning frameworks based on TrixInstruct. (1) w/o Planning: We split each instance in TrixInstruct into planning and reasoning parts. We train our model with only the reasoning part of the training instances. This can be considered as distilling the reasoning pattern of TabSQLify Nahid and Rafiei (2024). (2) w/o Reasoning: Similar to (1), we finetune the model with only the planning part of the training instances. (3) w/o Planning and Reasoning: We finetune the model to generate answers directly. This is similar to the end-to-end QA paradigm.
@@ 4.1 Ablation study
4.2 Prompting Efficiency Analysis
In Table 4, we compare the number of annotated in-context demonstrations and average API calls of different prompting methods on WikiTQ. It shows that our Plan-then-Reason framework can effectively reach the final answer with less in-context demonstrations and API calls. Binder Cheng et al. (2022b), Dater Ye et al. (2023), Chain-of-Table Wang et al. (2024) and TabSQLify Nahid and Rafiei (2024) require more than 10 annotated in-context examples for inference while our Plan-then-Reason framework only needs one in-context demonstration. Binder, Dater and MIX SC predict the answer with the self-consistency mechanism which needs at least 10 API calls. Chain-of-Table follows the style of ReAct Yao et al. (2023) with averagely 10.1 API calls to reach the final answer, while Plan-then-Reason only requires 2 calls on average.
@@ 4.2 Prompting Efficiency Analysis
4.3 Error Analysis
@@
We perform error analysis on the results of Plan-then-Reason with GPT-3.5-turbo and ProTrix on WikiTQ and FEVEROUS to analyze the potential weakness of our model. We categorize the error cases into the following types: (1) Planning error: the response fails to generate a correct plan to answer the question, (2) SQL error: The response generates SQL containing columns that do not exist at all or the SQL simply can not be executed. (3) Execution error: the execution result given by GPT-4 is wrong. (4) Reasoning error: the model achieves a correct answer with wrong reasoning. (5) False Negative: the gold answer is wrong or the executor misjudges the result. We show examples for each error type in Table 9 in Appendix. We also perform qualitative analysis shown in Appendix E.
Error Type WikiTQ FEVEROUS
ProTrix GPT-3.5 ProTrix GPT-3.5
Planning Error 22% 18% 36% 46%
SQL Error 32% 26% 22% 12%
Execution Error 30% 28% 10% 8%
Reasoning Error 12% 14% 26% 26%
False Negative 4% 14% 6% 8%
Table 5: Error analysis of ProTrix and GPT-3.5-turbo on WikiTQ and FEVEROUS.
Despite the effectiveness of our framework in both in-context learning and finetuning methods, we identify two primary sources of errors. The first error source is generating SQL queries for complex table content or program structures. The main challenge for WikiTQ is to write the correct SQL to perform complex operations on the table. The model sometimes generates SQLs with invalid program structures for questions that need multi-step table operations such as comparing the results of two maximum values under different conditions. The model also struggles to write SQL to process complex context such as isolating country names from table cells as John (ESP) where the data is embedded within the text. We think it is due to the limited coding ability in generating SQLs for multi-step operations or complex cell content. We plan to enhance the specific coding ability for SQL of ProTrix in future work.
The second main error source is decomposing complex queries into sub-questions. FEVEROUS challenges the model to decompose a claim into several sub-questions, digest the relevant table or sentences to answer each sub-question, and then combine the results to form the final answer. In some cases, ProTrix is observed to inaccurately assign sub-questions that are directly related to the sentences to be answered using tables, or it may overlook certain sub-questions during the decomposition of the original claim. We suspect that the error source of missing sub-questions is that the model can not retain all the sub-questions from the claim during planning due to its autoregressive generation nature. We plan to implement a memory mechanism of sub-questions to address this issue.
5 Related Work
Prompting Methods for LLMs
Large language models can be guided to solve tasks in a step-by-step manner Wei et al. (2022); Hao et al. (2023). Chen (2023) first utilizes Chain-of-Thought Wei et al. (2022) to enhance the reasoning of LLMs on tabular tasks and points out that textual reasoning can not generalize to large tables directly. Researchers prompt the model to select relevant rows and columns as one step in the chain of reasoning to enable LLM to focus on the following reasoning step Jiang et al. (2023); Ye et al. (2023); Wang et al. (2024). Chen et al. (2022) proposes Program-of-Thoughts (PoT) that answers a question in programming language. Compared with textual reasoning, program-based reasoning is executed by a code interpreter, achieving high-precision reasoning in complex tabular or mathematical questions. Binder Cheng et al. (2022b) binds LLMs as API calls within a Python or SQL program to address the program-unsolvable aspect of the queries. Liu et al. (2023b) proposes mix self-consistency that combines the potential of both textual and program-based reasoning. Researchers have also attempted ReAct Yao et al. (2023) style prompting for tabular tasks. ReAcTable Zhang et al. (2023b) reasons with SQL or Python tools in multiple turns. Chain-of-Table Wang et al. (2024) formats table reasoning as specific table operations. ReAct-style prompting mainly focuses on the next acting step based on feedback from the last action without considering the whole reasoning chain. Our Plan-then-Reason framework can plan the whole reasoning chain before acting and answer questions more efficiently and accurately (see efficiency analysis in Appendix 4.2) while blending the advantages of textual and program-based reasoning methods.
Finetuned Models
Various pre-trained models are proposed for tabular tasks Yin et al. (2020); Wang et al. (2021); Iida et al. (2021); Deng et al. (2022); Yang et al. (2022); Jiang et al. (2022); Liu et al. (2021). But they often are limited to one specific downstream finetuning task. As for models with generalizability, Liu et al. (2023a) mix symbolic SQL execution task with FLAN task to further finetune FLAN-T5 to improve zero-shot tabular question answering performance. Li et al. (2023) finetunes models with a large synthesized dataset of table manipulation and data augmentation to serve as a table-foundation model that understands table structures. TableLlama Zhang et al. (2023a) and StructLM Zhuang et al. (2024) collect an instructing tuning set that covers diverse tables and tasks and finetune Llama to obtain a generalist model without table pretraining. Compared with existing generalist models that are expected to generate answers directly, ProTrix is interpretable by generating accurate and faithful explanations.
6 Conclusions
In this paper, we propose a novel Plan-then-Reason framework to answer user queries over tables with sentence context, which analyzes the commonsense and concepts in the query and plans the reasoning steps over programs and natural languages. We construct an instruction tuning set TrixInstruct to finetune models to obtain such planning and reasoning abilities with only 6k examples. Experiments show that our resulting models ProTrix and ProTrix-Coder can generalize to unseen tabular tasks with sentence context and produce accurate and faithful explanations. Our work highlights the required abilities for generalist models over tabular tasks with sentence context, and paves the way for future research directions.
Limitations
Acknowledgement
This work is supported in part by NSFC (62161160339) and Beijing Science and Technology Program (Z231100007423011). We would like to thank the anonymous reviewers for their comments and suggestions. We also thank Rami Aly and Md Mahadi Hasan Nahid for the helpful discussions. For any correspondence, please contact Yansong Feng.
Appendix A Implementation Details
We fully finetune Llama-2 7B Touvron et al. (2023) and CodeLlama-7B Roziere et al. (2023) with our instruction tuning set with the context of length 4096. We set the learning rate as 5e-6 and the batch size as 32. The training process uses a cosine scheduler with a 3% period for 3 epochs. We utilize DeepSpeed training with ZeRO-3 stage Rasley et al. (2020). Our model is trained with 4 NVIDIA A40 GPUs (48GB) and the whole training process takes about 5 hours.
For finetuned SOTA methods in Table 1, we report the performance of OmniTab Jiang et al. (2022) for WikiTQ, TAPEX Liu et al. (2021) for WikiSQL, PASTA Gu et al. (2022) for TabFact, S3HQA Lei et al. (2023) for HybridQA and APOLLO Sun et al. (2022) for TATQA. For FEVEROUS, we run DCUF Hu et al. (2022) on our training and development set of FEVEROUS and obtain an accuracy of 75.9%. Notice that S3HQA uses a more precise sentence retriever compared to ours and DCUF leverages an additional retriever to select 25 table cells as input.
Appendix B Training Dataset Analysis
Our instruction tuning dataset is extracted from GPT-4 responses. We filter out the responses that have inconsistent final answers with the original dataset annotations, but the reasoning process of the responses in the training set has not been fully validated. We perform a quality analysis of our training dataset following the error types defined in § 4.3 We sample 50 cases with sentence context and 50 cases without sentence context from TrixInstruct to perform manual evaluation.
@@ Appendix B Training Dataset Analysis
Appendix C Analysis of Breakdown Performance
C.1 Program-Unsolvable Queries
@@
From Table 7, we can observe that ProTrix-Coder achieves the highest accuracy on program-unsolvable queries compared with finetuned methods. It suggests TrixInstruct can teach a model to understand commonsense and general concepts in the query and adaptatively plan to reason with programs or languages. ProTrix-Coder still falls behind TAPEX and TaCube on the program-solvable subset. But these models require table pretraining which is computationally expensive. ProTrix-Coder surpasses the previous generalist model by 1.5% and 4.7% on program unsolvable and solvable subsets, indicating the effectiveness of the proposed Plan-then-Reason framework.
@@
C.2 Combining Tables and Sentences
@@
Appendix D More Results
@@
Appendix E Case Study